West Highschool has 250 fewer students than twice the amount of East Highschool. Together they have 2858 students, how many students does each school have?
Answer by robertb(5830)

let w = #students in West, e = # students in East.
Then from the given,
w = 2e -250. Also, w +e = 2858.
from the 2nd equation, w = 2858 - e. Substituting into the 1st equation,
2858 -e = 2e - 250.
3108 = 2e +e, after transposition.
3108 = 3e,
1036 = e, and w = 2858-1036 = 1822.